OpenStudy (anonymous):

solve for x radical x+9-2 = 2

OpenStudy (saifoo.khan):

\[\sqrt{x+9}-2=2\]that?

OpenStudy (anonymous):

yup yup

OpenStudy (saifoo.khan):

x+9 = 4^2 x + 9 = 16 x = 16-9 x = 7

OpenStudy (anonymous):

ty :D

OpenStudy (saifoo.khan):

uw. :D

OpenStudy (lalaly):

\[\large{SAIFO}\]

OpenStudy (saifoo.khan):

\[\Huge LuLu.\]

OpenStudy (anonymous):

lozl thts big

OpenStudy (lalaly):

hehehe :D:D

OpenStudy (saifoo.khan):

;]